OpenStudy (bohotness):

1. In one month, Jillian made 36 local phone calls and 20 long distance calls. What was her ratio of local calls to long distance calls for that month? A. 9 : 5 B. 3 : 2 C. 2 : 1 D. 9 : 4

OpenStudy (bohotness):

@sleepyhead314

OpenStudy (anonymous):

36/20 = 9/5 = 9:5

OpenStudy (anonymous):

So the Answer would be?

OpenStudy (bohotness):

a

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

You got it!

OpenStudy (bohotness):

Ellie had 44 books in her bookcase. She counted 32 fiction books and the rest were nonfiction. What was the ratio of fiction books to nonfiction books in Ellie's bookcase? A. 8 to 3 B. 3 to 8 C. 8 to 11 D. 11 to 3

OpenStudy (bohotness):

@AnswerMyQuestions @AleshGames

OpenStudy (anonymous):

Subtract 44 - 32

OpenStudy (bohotness):

a?

OpenStudy (anonymous):

yes it is a

OpenStudy (bohotness):

44-32=12

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So simplify 32:12

OpenStudy (anonymous):

Divide them both by 4.

OpenStudy (anonymous):

32 \(\div\) 4 12 \(\div\) 4

OpenStudy (anonymous):

Then you put them up on the ratio.

OpenStudy (bohotness):

32\4=8 12\4=3

OpenStudy (bohotness):

so a